An engine has an inertia moment of 2 kgm². What moment of force must be exerted on it, so that the rotation frequency increases from 150 rpm to 400 rpm, during 5 s.


1
Expert's answer
2021-04-01T17:03:33-0400
"\tau=I\alpha\\\tau=(2)\frac{2\pi}{60}\frac{400-150}{5}=10.5\ Nm"